# Adaptive Gradient Normalization and Independent Sampling for (Stochastic) Generalized-Smooth Optimization

### v3

@@ 2 Related Work
A number of works also investigated the performance of diverse stochastic algorithms—extending beyond SGD and clipped SGD under the generalized-smooth condition. Wang et al. (2023); Faw et al. (2023); Hong & Lin (2024) studied AdaGrad (Duchi et al., 2011b) under generalized-smooth and affine variance assumption with different learning rate schemes. They all attain $\mathcal{\tilde{O}}(1/\sqrt{T})$ convergence rate under mild conditions. Xie et al. (2024a) studied trust-region methods convergence under generalized-smoothness. For stochastic acceleration methods under the generalized-smoothness condition, Zhang et al. (2020a) proposed a general clipping framework by leveraging momentum clipping and they achieve $\mathcal{O}(\epsilon^{-4})$ sample complexity under almost sure bounded noise assumption; Jin et al. (2021) studied normalized SGD with momentum under parameter-dependent learning rates schemes, which achieves $\mathcal{O}(\epsilon^{-4})$ sample complexity under generalized-smooth and affine variance assumption; Hübler et al. (2024) studied normalized SGD with momentum (Cutkosky & Mehta, 2020) under parameter-agnostic learning rates schemes, which establishes $\tilde{\mathcal{O}}({\epsilon^{-4}})$ convergence rate under bounded variance assumption. Chen et al. (2023); Reisizadeh et al. (2023) demonstrated that the SPIDER algorithm (Fang et al., 2018) can reach the optimal $\mathcal{O}(\epsilon^{-3})$ sample complexity under affine variance assumption for sum-type functions by employing a large batch size. Furthermore, Zhang et al. (2024b); Wang et al. (2024a; b); Li et al. (2023) explored the convergence of RMSprop (Hinton et al., 2012) and Adam (Kingma, 2014) under different noise assumptions, including coordinate-wise affine variance, expected affine variance, bounded variance and sub-gaussian distribution. The acceleration of sign-SGD was explored by Jiang et al. (2024); Sun et al. (2023b), and its generalized-smooth variants were subsequently analyzed by Crawshaw et al. (2022); Sun et al. (2023a), where they both obtain $\mathcal{O}(\epsilon^{-4})$ under almost-sure bound and bounded variance assumptions, respectively.
@@
Table 1: Comparison of existing clipping-based algorithms. Denote the stochastic gradient and clipping threshold as $\nabla f_{\xi}(w_{t})$ and $c$, respectively. Explanation of abbreviations: (GS) refers to the generalized-smoothness condition, and (LS) refers to the standard $L$-smoothness condition. (a.s. bound) means that for every sample, the stochastic gradient bias is almost surely bounded, i.e., $\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|\leq\tau_{2}$; (var bound) denotes the bounded variance condition, $\mathbb{E}_{\xi}[\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|^{2}]\leq\tau_{2}^{2}$; (a.s. affine) implies that for every sample, the stochastic gradient bias satisfies the affine bound $\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|\leq\tau_{1}\|\nabla F(w_{t})\|+\tau_{2}$ almost surely, for some $10\leq\tau_{1}<$ 1 and $\tau_{2}>0$. The term Clip-SGD refers to algorithms that use stochastic gradient clipping of the form $w_{t+1}=w_{t}-\gamma\min\{1,\frac{c}{\|\nabla f_{\xi}(w_{t})\|}\}\nabla f_{\xi}(w_{t})$; SGD refers to stochastic gradient descent; General-Clip denotes algorithms that incorporate momentum during clipping and updates $w_{t+1}=w_{t}-(\nu\min\{c,\frac{\gamma}{\|m_{t}\|}\}m_{t}+(1-\nu)\min\{c,\frac{\gamma}{\|\nabla f_{\xi}(w_{t})\|}\}\nabla f_{\xi}(w_{t}))$, where $m_{t+1}=(1-\vartheta)m_{t}+\vartheta\nabla f_{\xi}(w_{t})$. The label $\epsilon$-stationary indicates that the algorithm can converge to an arbitrary target error. For Koloskova et al. (2023), we report their convergence result under learning rate $\gamma=\epsilon^{2}/\left(\tau_{2}^{2}(L_{0}+cL_{1})\right)$ with $c=\tau_{2}/\epsilon$, where their analysis guarantees convergence to an $\epsilon$-stationary point.
@@ 3 Notations
[…]
@@ 5 Adaptive Gradient Normalization for Deterministic Generalized-Smooth Optimization
[…]
@@ 6.1 Normalized SGD and Its Limitations
[…]
@@ 6.2 Independently-and-Adaptively Normalized SGD
[…]
where $\xi$ and $\xi^{\prime}$ are samples draw independently from the underlying data distribution. Intuitively, the independence between $\xi$ and $\xi^{\prime}$ decorrelates the denominator from the numerator, making it an unbiased stochastic gradient estimator (up to a scaling factor). Specifically, we have that
@@
Moreover, as we show later under mild assumptions, the scaling factor $\mathbb{E}\big{[}\|\nabla f_{\xi^{\prime}}(w)\|^{-\beta}\big{]}$ can be roughly bounded by the full gradient norm and hence resembling the full-batch AN-GD update. Based on this idea, we formally propose the following independently-and-adaptively normalized SGD (IAN-SGD) algorithm (see equation 14), where $A$, $\Gamma$,$\delta$ are positive constants, $\nabla f_{\xi}(w_{t})$, $\nabla f_{\xi^{\prime}}(w_{t})$ corresponds to the stochastic gradient associated with two independent samples $\xi$, $\xi^{\prime}$ sampled at time $t$.
[…]
@@ 6.3 Convergence Analysis of IAN-SGD
There exists $00\leq\tau_{1}<1,\tau_{2}>$ 0 such that for any $w\in\mathbf{R}^{d}$,
@@ 6.3.1 Proof outline and novelty
$\frac{\gamma}{4h_{t}^{\beta}}\big{\|}\nabla F(w_{t})\big{\|}^{2}\leq$ $\mathbb{E}_{\xi}\big{[}F(w_{t})-F(w_{t+1})\big{]}+\frac{1}{2}(L_{0}+L_{1})\gamma^{2}(1+4\tau_{2}^{2})^{2}.$
@@
$\mathbb{E}_{\xi}\frac{\|\nabla f_{\xi}(w_{t})\|^{2}}{h_{t}^{2\beta}}=\mathbb{E}_{\xi}\Big{[}\Big{(}\|\nabla F(w_{t})\|^{2}+\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|^{2}+2\langle\nabla F(w_{t}),\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\rangle\Big{)}/{h_{t}^{2\beta}}\Big{]}.$
@@ 7 Experiments
We conduct numerical experiments to compare IAN-SGD with other state-of-the-art stochastic algorithms, including the standard SGD (Ghadimi & Lan, 2013), normalized SGD (NSGD), clipped SGD (Zhang et al., 2019), SPIDER (Fang et al., 2018), normalized SGD with momentum (Cutkosky & Mehta, 2020)(NSGDm) etc. Compared with previous works that study clipped SGD (Zhang et al., 2019; 2020a; 2020b; Koloskova et al., 2023; Reisizadeh et al., 2023), the main change of IAN-SGD lies in its usage of independent sampling and adaptive normalization. To empirically validate the effect of adaptive normalization in accordance with Theorem 1, our experiments focus on nonconvex phase retrieval, distributionally robust optimization (DRO), and deep neural network training. Prior works (Zhou et al., 2016; Jin et al., 2021; Liu et al., 2022a; Scaman et al., 2022) have shown that phase retrieval and deep neural networks satisfy generalized PŁ conditions under mild assumptions. Accordingly, we evaluate the convergence of all algorithms in terms of total sample complexity (for phase retrieval and DRO) and number of epochs (for deep neural networks). Furthermore, to align the DRO formulation with our theoretical assumptions, we focus on regression tasks and adopt the $\chi^{2}$-divergence to model distributional uncertainty.
@@ 7.1 Nonconvex Phase Retrieval
Such nonconvex function satisfies generalized-smooth with parameter $\alpha=\frac{2}{3}$. In this experiment, we generate the initialization $w_{0}\sim\mathcal{N}(1,6)$ and the underlying true object $w^{*}\sim\mathcal{N}(0,0.5)$ with dimension $d=100$. We take $m=3$k measurements with $a_{r}\sim\mathcal{N}(0,0.5)$ and $n_{r}\sim\mathcal{N}(0,16)$. We implement all of the stochastic algorithms in original form as described in previous literature. We set batch size $|B|=64$, and for IAN-SGD, we choose a small independent batch size $|B^{\prime}|=4$. We fine-tuned learning rate for all algorithms, i.e., $\gamma=5e^{-5}$ for SGD, $\gamma=0.2$ for NSGD and NSGD with momentum, $\gamma=0.6$ for clipped SGD, and $\gamma=0.25$ for SPIDER and IAN-SGD. We set the maximal gradient clipping constant as $20$, $\delta=1e^{-3}$ for both clipped SGD and IAN-SGD. And we set normalization parameter $\beta=\frac{2}{3}$. Figure 2 (left) shows the comparison of objective value versus sample complexity. It can be observed that IAN-SGD consistently converges faster than other baseline algorithms.
@@ 7.3 Deep Neural Networks
For SGD, Adam and Adagrad, we utilize PyTorch built-in optimizer (Paszke et al., 2019) to implement training pipelines. We implement training pipelines for NSGD, NSGD with momentum, clipped SGD and IAN-SGD. The normalization constant is computed through all model parameters at each iteration. The detailed algorithm settings are as follows. For batch size, all algorithms use $B=128$, and $B^{\prime}=32$ for IAN-SGD. For stochastic momentum moving average parameter, we set it as $0.9$, $0.99$ for Adam, and 0.25 for normalized SGD with momentum. For clipping threshold used in clipped SGD and IAN-SGD, we set them as 2 and $\delta=1e^{-1}$. The normalization power used for IAN-SGD is $\beta=\frac{2}{3}$. We fine-tuned learning rate for all algorithms, i.e., $\gamma=1e^{-3}$ for SGD, Adam and Adagrad, $\gamma=1e^{-1}$ for NSGD and NSGD with momentum, $\gamma=2e^{-1}$ for clipped SGD and IAN-SGD. We trained ResNet18, ResNet50 on CIFAR10 dataset for 30 epochs and plot the training loss in Figure 3. Figure 3 (left) shows the training loss of ResNet18, Figure3 (right) shows the training loss of ResNet50. As we can see from these figures, the (pink) loss curve optimized by IAN-SGD indicates fast convergence rate comparable with several baselines, including SGD, NSGD NSGDm, clipped SGD, which demonstrate the effectiveness of IAN-SGD framework.
@@ Appendix C Proof for Nonconvex Phase Retrieval
=$\overset{(i)}{=}\Big{|}\big{|}a_{\xi}^{\top}w\big{|}^{3}-y_{\xi}\big{|}a_{\xi}^{\top}w\big{|}\Big{|}^{\frac{2}{3}}\big{\|}a_{\xi}\big{\|}^{\frac{2}{3}}$
@@ Appendix D Proof of Descent Lemma under Generalized PŁ condition
=$\overset{(ii)}{=}$ $-\gamma\big{\|}\nabla f\big{(}w_{t}\big{)}\big{\|}^{2-\beta}+\frac{\gamma}{4}\big{(}2L_{0}\gamma\cdot\big{\|}\nabla f\big{(}w_{t}\big{)}\big{\|}^{2-2\beta}+2L_{1}\gamma\cdot\big{\|}\nabla f\big{(}w_{t}\big{)}\big{\|}^{2+\alpha-2\beta}\big{)}$
@@
where (i) follows from the descent lemma 1; (ii) follows from the AN-GD update rule by replacing $w_{t+1}-w_{t}$ with $-\gamma\frac{\nabla f(w_{t})}{\|\nabla f(w_{t})\|^{\beta}}$; (iii) follows from the technical lemma 5 above proved in Chen et al. (2023) by letting $\omega^{\prime}=2-\beta$, $\Delta=\beta$ and substituting equation 29 to $2L_{0}\gamma\|\nabla f(w_{t})\|^{2-2\beta}$ and $2L_{1}\gamma\|\nabla f(w_{t})\|^{2+\alpha-2\beta}$ (Let $x=\|\nabla f(w_{t})\|$ and $C=2L_{0}\gamma\text{ or }2L_{1}\gamma$ , $C\in[0,1]$ because $\epsilon\leq 1/2\mu$); (iv) follows from the fact that $a^{\tau}+b^{\tau}\leq(a+b)^{\tau}$ with $\tau={2}/{\beta}-1\geq 1$ and $a,b\geq 0$, $(v)$ follows from the step size rule $\gamma={(2\mu\epsilon)^{\beta/\rho}}/{(8(L_{0}+L_{1})+1)}$; (vi) follows from the facts that $12L_{0}+2L_{1}\leq 8(L_{0}+L_{1})+$ 1 and $(2L_{0}+2L_{1})^{\frac{2}{\beta}-1}\leq(8(L_{0}+L_{1})+1)^{\frac{2}{\beta}-1}$ when $\beta>0$.
@@
=$\overset{(ii)}{=}$ $-\gamma\big{\|}\nabla f\big{(}w_{t}\big{)}\big{\|}^{2}+\frac{\gamma}{4}\big{(}2L_{0}\gamma\cdot\big{\|}\nabla f\big{(}w_{t}\big{)}\big{\|}^{2}+2L_{1}\gamma\cdot\big{\|}\nabla f\big{(}w_{t}\big{)}\big{\|}^{2}\big{)}$
@@
where (i) follows from the descent lemma 1 with $\alpha=0$; (ii) follows from the update rule of AN-GD by replacing $w_{t+1}-w_{t}$ with $-\gamma\nabla f(w_{t})$ (Normalization term vanishes due to $\beta=0$); (iii) follows the fact that $12L_{0}+2L_{1}\leq 8(L_{0}+L_{1})+$ 1; (iv) follows from the fact $\frac{\gamma}{4}\leq\frac{\gamma}{2}$ and (v) holds by adding additional positive number $\frac{\gamma}{4}(2\mu\epsilon)^{\frac{2}{\rho}}$.

### v4

@@ 2 Related Work
A number of works also investigated the performance of diverse stochastic algorithms—extending beyond SGD and clipped SGD under the generalized-smooth condition. Wang et al. (2023); Faw et al. (2023); Hong & Lin (2024) studied AdaGrad (Duchi et al., 2011b) under generalized-smooth and affine variance assumption with different learning rate schemes. They all attain $\mathcal{\tilde{O}}(1/\sqrt{T})$ convergence rate under mild conditions. Xie et al. (2024a) studied trust-region methods convergence under generalized-smoothness. For stochastic acceleration methods under the generalized-smoothness condition, Zhang et al. (2020a) proposed a general clipping framework by leveraging momentum clipping and they achieve $\mathcal{O}(\epsilon^{-4})$ sample complexity under almost sure bounded noise assumption; Jin et al. (2021) studied normalized SGD with momentum under parameter-dependent learning rates schemes, which achieves $\mathcal{O}(\epsilon^{-4})$ sample complexity under generalized-smooth and affine variance assumption; Hübler et al. (2024) studied normalized SGD with momentum (Cutkosky & Mehta, 2020) under parameter-agnostic learning rates schemes, which establishes $\tilde{\mathcal{O}}({\epsilon^{-4}})$ convergence rate under bounded variance assumption. Chen et al. (2023); Reisizadeh et al. (2023) demonstrated that the SPIDER algorithm (Fang et al., 2018) can reach the optimal $\mathcal{O}(\epsilon^{-3})$ sample complexity under affine variance assumption for sum-type functions by employing a large batch size. Furthermore, Zhang et al. (2024b); Wang et al. (2024a; b); Li et al. (2023) explored the convergence of RMSprop (Hinton et al., 2012) and Adam (Kingma, 2014) under different noise assumptions, including coordinate-wise affine variance, expected affine variance, bounded variance and sub-gaussian distribution. The acceleration of sign‑SGD was explored by Jiang et al. (2024); Sun et al. (2023b), and its generalized‑smooth variants were subsequently analyzed by Crawshaw et al. (2022); Sun et al. (2023a), where they both obtain $\mathcal{O}(\epsilon^{-4})$ under almost‑sure bound and bounded variance assumptions, respectively.
@@
Table 1: Comparison of existing clipping-based algorithms. Denote the stochastic gradient and clipping threshold as $\nabla f_{\xi}(w_{t})$ and $c$, respectively. Explanation of abbreviations: (GS) refers to the generalized-smoothness condition, and (LS) refers to the standard $L$-smoothness condition. (a.s. bound) means that for every sample, the stochastic gradient bias is almost surely bounded, i.e., $\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|\leq\tau_{2}$; (var bound) denotes the bounded variance condition, i.e., $\mathbb{E}_{\xi}[\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|^{2}]\leq\tau_{2}^{2}$; (a.s. affine) implies that for every sample, the stochastic gradient bias satisfies the affine bound $\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|\leq\tau_{1}\|\nabla F(w_{t})\|+\tau_{2}$ almost surely, for some $0\leq\tau_{1}<1$ and $\tau_{2}>0$. The term Clip-SGD refers to algorithms that use stochastic gradient clipping of the form $w_{t+1}=w_{t}-\gamma\min\{1,\frac{c}{\|\nabla f_{\xi}(w_{t})\|}\}\nabla f_{\xi}(w_{t})$; SGD refers to stochastic gradient descent; General-Clip denotes algorithms that incorporate momentum during clipping and updates $w_{t+1}=w_{t}-(\nu\min\{c,\frac{\gamma}{\|m_{t}\|}\}m_{t}+(1-\nu)\min\{c,\frac{\gamma}{\|\nabla f_{\xi}(w_{t})\|}\}\nabla f_{\xi}(w_{t}))$, where $m_{t+1}=(1-\vartheta)m_{t}+\vartheta\nabla f_{\xi}(w_{t})$. The label $\epsilon$-stationary indicates that the algorithm can converge to an arbitrary target error. For Koloskova et al. (2023), we report their convergence result under learning rate $\gamma=\epsilon^{2}/\left(\tau_{2}^{2}(L_{0}+cL_{1})\right)$ with $c=\tau_{2}/\epsilon$, where their analysis guarantees convergence to an $\epsilon$-stationary point.
@@ 3 Notations
Throughout the work, we denote $D$ as a set of training examples, $\xi$ as an example sampled from the training dataset $D$, $f(\cdot)$ as the objective function, $f_{\xi}(\cdot)$ as the objective function associated with data sample $\xi$, $F(w)=\mathbb{E}_{\xi\sim\mathbb{P}}[f_{\xi}(w)]$ as the expected loss over all samples following distribution $\mathbb{P}$, and $\|\cdot\|$ as $\ell_{2}$-norm over Euclidean space. Let $w\in\mathbf{R}^{d}$ denote the parameters of function $f$ (i.e., the weights for linear model and neural networks), $f^{*}$ denote the infimum of $f$. $L_{0},L_{1}$ and $\alpha$ are positive constants used to characterize the generalized-smoothness condition (see Assumption 1). $\rho,\mu$ are positive constant used to describe generalized PŁ-condition (see Assumption 2). $\tau_{1},\tau_{2}$ characterize the stochastic gradient noise (see Assumption 4). $\gamma,\beta$ denote the learning rate and normalization power used in adaptive normalization, respectively. In the Distributionally Robust Optimization (DRO) experiments (see Section 7), we define $L(w,\eta)$ with parameter $w$ and dual variable $\eta$, as the dual objective derived from a regularized $\phi$-divergence primal DRO problem.
@@ 5 Adaptive Gradient Normalization for Deterministic Generalized-Smooth Optimization
$\text{(AN-GD)}\quad w_{t+1}=w_{t}-\gamma\frac{\nabla f(w_{t})}{\|\nabla f(w_{t})\|^{\beta}},$ (6)
@@ 6.1 Normalized SGD and Its Limitations
$\text{(AN-SGD)}\quad w_{t+1}=w_{t}-\gamma\frac{\nabla f_{\xi}(w_{t})}{\|\nabla f_{\xi}(w_{t})\|^{\beta}}.$ (11)
@@ 6.2 Independently-and-Adaptively Normalized SGD
$\text{(IAN-SG estimator)}\quad\frac{\nabla f_{\xi}(w)}{\|\nabla f_{\xi^{\prime}}(w)\|^{\beta}},$ (12)
where $\xi$ and $\xi^{\prime}$ are samples drawn independently from the underlying data distribution. Intuitively, the independence between $\xi$ and $\xi^{\prime}$ decorrelates the denominator from the numerator, making it an unbiased stochastic gradient estimator (up to a scaling factor). Specifically, we have that
@@
Moreover, as we show later under mild assumptions, the scaling factor $\mathbb{E}\big[\|\nabla f_{\xi^{\prime}}(w)\|^{-\beta}\big]$ can be roughly bounded by the full gradient norm and hence resembling the full-batch AN-GD update. Based on this idea, we formally propose the following independently-and-adaptively normalized SGD (IAN-SGD) algorithm (see equation 14), where $A$, $\Gamma$$\delta$ are positive constants, $\nabla f_{\xi}(w_{t})$, $\nabla f_{\xi^{\prime}}(w_{t})$ corresponds to the stochastic gradient associated with two independent samples $\xi$, $\xi^{\prime}$ sampled at time $t$.
$\text{(IAN-SGD):}~~w_{t+1}=w_{t}-\gamma\frac{\nabla f_{\xi}(w_{t})}{h_{t}^{\beta}},$~~
@@ 6.3 Convergence Analysis of IAN-SGD
There exists $0\leq\tau_{1}<1,\tau_{2}>0$ such that for any $w\in\mathbf{R}^{d}$,
@@ 6.3.1 Proof outline and novelty
$\mathbb{E}_{\xi^{\prime}}\big[\frac{\gamma}{4h_{t}^{\beta}}\big\|\nabla F(w_{t})\big\|^{2}\big]\leq$ $\mathbb{E}_{\xi,\xi^{\prime}}\big[F(w_{t})-F(w_{t+1})\big]+\frac{1}{2}(L_{0}+L_{1})\gamma^{2}(1+4\tau_{2}^{2})^{2}.$
@@
$\mathbb{E}_{\xi}\Big[\frac{\|\nabla f_{\xi}(w_{t})\|^{2}}{h_{t}^{2\beta}}\Big]=\mathbb{E}_{\xi}\Big[\Big(\|\nabla F(w_{t})\|^{2}+\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|^{2}+2\langle\nabla F(w_{t}),\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\rangle\Big)/{h_{t}^{2\beta}}\Big].$
@@ 7 Experiments
We conduct numerical experiments to compare IAN-SGD with other state-of-the-art stochastic algorithms, including the standard SGD (Ghadimi & Lan, 2013), normalized SGD (NSGD), clipped SGD (Zhang et al., 2019), SPIDER (Fang et al., 2018), normalized SGD with momentum (Cutkosky & Mehta, 2020)(NSGDm) etc. Compared with previous works that study clipped SGD (Zhang et al., 2019; 2020a; 2020b; Koloskova et al., 2023; Reisizadeh et al., 2023), the main change of IAN-SGD lies in its usage of independent sampling and adaptive normalization. To empirically validate the effect of adaptive normalization in accordance with Theorem 1, our experiments focus on nonconvex phase retrieval, distributionally robust optimization (DRO), and deep neural network training. Prior works (Zhou et al., 2016; Jin et al., 2021; Liu et al., 2022a; Scaman et al., 2022) have shown that phase retrieval and deep neural networks satisfy generalized PŁ conditions under mild assumptions. Accordingly, we evaluate the convergence of all algorithms in terms of total sample complexity (for phase retrieval and DRO) and number of epochs (for deep neural networks). Furthermore, to align the DRO formulation with our theoretical assumptions, we focus on regression tasks and adopt the $\chi^{2}$‑divergence to model distributional uncertainty. All experiments were conducted on a PC computer equipped with a 2424‑core CPU, 3232GB of RAM, and a single NVIDIA RTX40904090 GPU, running Python 3.8.
@@ 7.1 Nonconvex Phase Retrieval
Such nonconvex function satisfies generalized-smooth with parameter $\alpha=\frac{2}{3}$. In this experiment, we generate the initialization $w_{0}\sim\mathcal{N}(1,6)$ and the underlying true object $w^{*}\sim\mathcal{N}(0,0.5)$ with dimension $d=100$. We take m=3m=3k measurements with $a_{r}\sim\mathcal{N}(0,0.5)$ and $n_{r}\sim\mathcal{N}(0,16)$. We implement all of the stochastic algorithms in original form as described in previous literature. We set batch size $|B|=64$, and for IAN-SGD, we choose a small independent batch size $|B^{\prime}|=4$. We fine-tuned learning rate for all algorithms, i.e., $\gamma=5e^{-5}$ for SGD, $\gamma=0.2$ for NSGD and NSGD with momentum, $\gamma=0.6$ for clipped SGD, and $\gamma=0.25$ for SPIDER and IAN-SGD. We set the maximal gradient clipping constant as $20$, $\delta=1e^{-3}$ for both clipped SGD and IAN-SGD. And we set normalization parameter $\beta=\frac{2}{3}$. Figure 2 (left) shows the comparison of objective value versus sample complexity. It can be observed that IAN-SGD consistently converges faster than other baseline algorithms.
@@ 7.3 Deep Neural Networks
For SGD, Adam and Adagrad, we utilize PyTorch built-in optimizer (Paszke et al., 2019) to implement training pipelines. We implement training pipelines for NSGD, NSGD with momentum, clipped SGD and IAN-SGD. The normalization constant is computed through all model parameters at each iteration. The detailed algorithm settings are as follows. For batch size, all algorithms use $B=128$, and $B^{\prime}=32$ for IAN-SGD. For stochastic momentum moving average parameter, we set it as $0.9$, $0.99$ for Adam, and 0.25 for normalized SGD with momentum. For clipping threshold used in clipped SGD and IAN-SGD, we set them as 2 and $\delta=1e^{-1}$. The normalization power used for IAN-SGD is $\beta=\frac{2}{3}$. We fine-tuned learning rate for all algorithms, i.e., $\gamma=1e^{-3}$ for SGD, Adam and Adagrad, $\gamma=1e^{-1}$ for NSGD and NSGD with momentum, $\gamma=2e^{-1}$ for clipped SGD and IAN-SGD. We trained ResNet18, ResNet50 on CIFAR10 dataset for $60,80$ epochs and plot the training loss in Figure 3. Figure 3 (left) shows the training loss of ResNet18, Figure3 (right) shows the training loss of ResNet50. As we can see from these figures, the (pink) loss curve optimized by IAN-SGD indicates fast convergence rate comparable with several baselines, including SGD, NSGD NSGDm, clipped SGD, which demonstrate the effectiveness of IAN-SGD framework.
@@ Appendix C Proof for Nonconvex Phase Retrieval
$\overset{(i)}{=}\Big|\big|a_{\xi}^{\top}w\big|^{3}-y_{\xi}\big|a_{\xi}^{\top}w\big|\Big|^{\frac{2}{3}}\big\|a_{\xi}\big\|^{\frac{2}{3}}$
@@ Appendix D Proof of Descent Lemma under Generalized PŁ condition
$\overset{(ii)}{=}$ $-\gamma\big\|\nabla f\big(w_{t}\big)\big\|^{2-\beta}+\frac{\gamma}{4}\big(2L_{0}\gamma\cdot\big\|\nabla f\big(w_{t}\big)\big\|^{2-2\beta}+2L_{1}\gamma\cdot\big\|\nabla f\big(w_{t}\big)\big\|^{2+\alpha-2\beta}\big)$
@@
where (i) follows from the descent lemma 1; (ii) follows from the AN-GD update rule by replacing $w_{t+1}-w_{t}$ with $\gamma\frac{\nabla f(w_{t})}{\|\nabla f(w_{t})\|^{\beta}}$; (iii) follows from the technical lemma 5 above proved in Chen et al. (2023) by letting $\omega^{\prime}=2-\beta$, $\Delta=\beta$ and substituting equation 29 to $2L_{0}\gamma\|\nabla f(w_{t})\|^{2-2\beta}$ and $2L_{1}\gamma\|\nabla f(w_{t})\|^{2+\alpha-2\beta}$ (Let $x=\|\nabla f(w_{t})\|$ and $C=2L_{0}\gamma\text{ or }2L_{1}\gamma$ , $C\in[0,1]$ because $\epsilon\leq 1/2\mu$); (iv) follows from the fact that $a^{\tau}+b^{\tau}\leq(a+b)^{\tau}$ with $\tau={2}/{\beta}-1\geq 1$ and $a,b\geq 0$, $(v)$ follows from the step size rule $\gamma={(2\mu\epsilon)^{\beta/\rho}}/{(8(L_{0}+L_{1})+1)}$; (vi) follows from the facts that $2L_{0}+2L_{1}\leq 8(L_{0}+L_{1})+1$ and $(2L_{0}+2L_{1})^{\frac{2}{\beta}-1}\leq(8(L_{0}+L_{1})+1)^{\frac{2}{\beta}-1}$ when $\beta>0$.
@@
$\overset{(ii)}{=}$ $-\gamma\big\|\nabla f\big(w_{t}\big)\big\|^{2}+\frac{\gamma}{4}\big(2L_{0}\gamma\cdot\big\|\nabla f\big(w_{t}\big)\big\|^{2}+2L_{1}\gamma\cdot\big\|\nabla f\big(w_{t}\big)\big\|^{2}\big)$
@@
where (i) follows from the descent lemma 1 with $\alpha=0$; (ii) follows from the update rule of AN-GD by replacing $w_{t+1}-w_{t}$ with $\gamma\nabla f(w_{t})$ (Normalization term vanishes due to $\beta=0$); (iii) follows the fact that $2L_{0}+2L_{1}\leq 8(L_{0}+L_{1})+1$; (iv) follows from the fact $\frac{\gamma}{4}\leq\frac{\gamma}{2}$ and (v) holds by adding additional positive number $\frac{\gamma}{4}(2\mu\epsilon)^{\frac{2}{\rho}}$.